OpenStudy (tiffany_rhodes):

Find the critical point of f(x,y)=x+y+7/xy. I know you find the critical point by setting the gradient equal to the zero vector and solving the system of equations for x and y.

OpenStudy (tiffany_rhodes):

are the partial derivatives df/dx = 1 -7y/(xy)^2 and df/dy = 1 -7x/(xy)^2

OpenStudy (tiffany_rhodes):

?

OpenStudy (johnweldon1993):

\[\large x + y + \frac{7}{xy}\] right?

OpenStudy (tiffany_rhodes):

yes

OpenStudy (johnweldon1993):

Went through the whole thing to say no...but realized you just wrote it different....but yes that is correct

OpenStudy (tiffany_rhodes):

Okay, so you set each of the partial derivatives equal to 0 and then solve for x and y correct?

OpenStudy (johnweldon1993):

Correct

OpenStudy (tiffany_rhodes):

to find the critical point(s) of the function?

OpenStudy (tiffany_rhodes):

Okay, so I'm having trouble with that. I was just curious about the partial derivatives to make sure I had the equations correct.

OpenStudy (phi):

1 -7y/(xy)^2 = 0 means 1 = 7/(x^2 y) x^2 y = 7 do the same for the other , df/dy

OpenStudy (johnweldon1993):

Ah...well as I wrote them out I have them as \[\large f_x = 1 - \frac{7}{x^2y}\] ad \[\large f_y = 1 - \frac{7}{xy^2}\] So lets do that 'x' out \[\large 1 = \frac{7}{x^2y}\] ahh phii beat me to it lol

OpenStudy (phi):

you should get x^2 y = 7 x y^2 = 7 divide these equations to get x/y= 7 x= 7y use that in the second equation 7 y * y^2 = 7 solve for y

OpenStudy (tiffany_rhodes):

Wow, I didn't realize the top y cancelled out. That makes things a lot simpler. I guess I've been out of algebra too long. Okay, so I got y = 7/x^2 when I solved the first equation

OpenStudy (phi):

yes. see the post above yours for how to continue

OpenStudy (tiffany_rhodes):

Thanks for the help!

OpenStudy (tiffany_rhodes):

both of you*